You are considering an investment in Justus Corporation's stock, which is expected to pay a dividend of $1.75 a share at the end of the year (D1 = $1.75) and has a beta of 0.9. The risk-free rate is 5.5%, and the market risk premium is 4.5%. Justus currently sells for $29.00 a share, and its dividend is expected to grow at some constant rate, g. The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the question below.

Assuming the market is in equilibrium, what does the market believe will be the stock price at the end of 3 years? (That is, what is ?) Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

We use the CAPM to find cost of equity i.e Rf + Risk Premium * Beta

= 5.5 + 4.5 * 0.9

= 9.55%

Using gordon growth model to find the growth

Price = Expected Dividend / (Ke - g)

29 = 1.75 / ( 9.55 - g)

6.03% = 9.55 % -g

g = 3.5155%

Expected dividend at end of 3 years = 1.75 (1 + 3.5155%)^3

= 1.94

Expected Share Price at end of 3 years = 1.94 / ( 9.55 - 3.5155)

= 32.15 or 32.17 (Round off Errors)